OpenStudy (anonymous):

which choice is the simple formula for the nth term of the following arithmetic sequence? 5,1, -3, -7, -11,... A. -4n+9 B. 4n+1 c. 4n+9 d. -4n+1 e. 5n+4

OpenStudy (jdoe0001):

what do you think is the difference from term to term? what makes the 2nd term 1, the 3rd term -3 and the 4th term -7?

OpenStudy (anonymous):

They are all being subracted by 4.

OpenStudy (anonymous):

**subtracted

OpenStudy (jdoe0001):

ahhh, well then, the \(\bf \text{ common ratio = -4}\) so use that in the sequence equation \(\bf \large \color{blue}{a_n = a_1 +(n-1)d}\\ a_n = n^{th} \text{ term}\\ a_1 = \text{ 1st term in the sequence}\\ d = \text{common ratio}\)

OpenStudy (anonymous):

that would make it -4n+9

OpenStudy (jdoe0001):

\(\large \bf a_n = 5 +(n-1)(-4) \implies a_n = 5 +(-4n+4)\)

OpenStudy (jdoe0001):

so simplifying that, will give us?

OpenStudy (anonymous):

-4n+9

OpenStudy (jdoe0001):

yes